OpenStudy (watermelon14):

if f(x)= sin2x+cosx, what is f(pi/4)

OpenStudy (watermelon14):



OpenStudy (watermelon14):

that is the answer but i don't know how.

OpenStudy (michele_laino):

it is simple: \[f\left( {\frac{\pi }{4}} \right) = \sin \left( {2 \cdot \frac{\pi }{4}} \right) + \cos \left( {\frac{\pi }{4}} \right) = ...?\]

OpenStudy (watermelon14):

oh wow thanks!